HUNTING DOWN THOSE BENDING VIBRATIONS OF LINEAR CARBON CLUSTERS

\underline{THOMAS F. GIESEN}, PETRA NEUBAUER-GUENTHER, GUIDO FUCHS, and GISBERT WIN-NEWISSER, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; ROBERT A. PROVENCAL, FRANK N. KEUTSCH, ALAN VAN ORDEN, and RICHARD J. SAYKALLY, Department of Chemistry, University of California, Berkeley, CA 94720, USA.

Pure carbon chain molecules, C_{n} (n = 3, 4, 5, ...) constitute an elusive but extremely important class of carbon clusters, not only from a spectroscopical point of view, but also from astro-chemical aspects, since many of the interstellar molecules contain carbon chains as backbones. Produced in supersonic jets, these highly reactive molecules can be characterised spectroscopically by their antisymmetric stretching modes in the IR-part of the electromagnetic spectrum. The carbon clusters can also be identified by their energetically low lying bending vibrations which occur in the terahertz region. These low ro-vibrational transitions of pure carbon chains and other centro-symmetric molecules which have no permanent dipole moment, provide a new way for the astrophysical detection of these species in cold interstellar clouds. Among the various carbon clusters studied, we have detected rather recently the existence of gas phase linear C_{10} and C_{8} by IR-tunable diode laser spectroscopy. These molecules show antisymmetric stretching modes at 2074 cm^{−1} and 2068 cm^{−1} respectively. The electronic ground state is a ^{3}Σ for even numbered and a ^{1}Σ for odd numbered carbon chains. The triplet splittings of rovibrational transitions of even numbered clusters were found to increase with the length of the chains. For C_{4} and C_{6} the splitting is barely to be seen but clearly resolved in the case of C_{8} and C_{10}.

C_{3} is so far the only carbon chain cluster which has been precisely investigated by its low bending ro-vibrational spectrum. We present our recent results on highly accurate measurements, which confirm our detection of C_{3} near the galactic center in Sgr B2 and data obtained from the ISO satellite recently.